OpenStudy (anonymous):

Find tan and sin;sec=4,sin>0 PLZ EXPLAIN

OpenStudy (ghazi):

sec can never be 4

OpenStudy (ghazi):

cos varies from -1 to +1

OpenStudy (anonymous):

it says in my text book

OpenStudy (ghazi):

then its wrong, maximum value of sec is 1 and minimum is -1

OpenStudy (anonymous):

using the pythagorean identities

OpenStudy (anonymous):

answer to this is in my text book hold on i will send u the answer

OpenStudy (ghazi):

then is not the sec theta =4 its the length of the side which is \[4^2=1^2+x^2-----x= \sqrt{16-1}=\sqrt{15}\]

OpenStudy (anonymous):